OpenStudy (anonymous):

Rationalize the denominator of 10 over quantity of 8 minus 5i

OpenStudy (anonymous):

multiply top and bottom by the conjuage of the denominator the conjugate of \(a+bi\) is \(a-bi\) and this works because \[(a+bi)(a-bi)=a^2+b^2\] a real number in your case \[\frac{10}{8-5i}=\frac{10}{8-5i}\times \frac{8+5i}{8+5i}\] \[=\frac{10(8-5i)}{8^2+5^2}\] simplify from there

OpenStudy (anonymous):

is it 80-50i over 69?

OpenStudy (anonymous):

i mean 89

OpenStudy (one098):

@sammixboo I got 80-50i/89 but that isnt one of the choices..

sammixboo (sammixboo):

Whoa... Old question :P

OpenStudy (one098):

lol yeah sorry

sammixboo (sammixboo):

Okay let me try in a few minutes. My mom just called me in to eat

OpenStudy (one098):

ok

OpenStudy (one098):

@Jhannybean

OpenStudy (jhannybean):

@satellite73 you cleared up my confusion about the technicalities involving conjugates of complex numbers, lol. thank you.

OpenStudy (one098):

but I'm still as confused as always! .-.

OpenStudy (jhannybean):

How @satellite73 solved it was to simplify the fraction with the imaginary number at the bottom by multiplying by the conjugate, that is the opposite of the function. When he did this, he got : \(8^2 +40i -40i -25i^2\) We know that \(i^2 = -1\) so substituting that into the function for i^2 he simpplified it to \(8^2 +5^2\)

OpenStudy (jhannybean):

All the OP has to do is expand the numerator.

OpenStudy (one098):

I got 80-50i/89 but again, that isn't one of the choices...

OpenStudy (jhannybean):

Then their problem might be wrong.

OpenStudy (one098):

:/ ok then